OpenStudy (anonymous):

evaluate: 2sinxcosx-cosx / 1-sinx+sin^2x-cos^2x

OpenStudy (anonymous):

\[2sinxcosx-cosx \over 1-sinx+\sin ^{2}x-\cos ^{2}x\]

OpenStudy (anonymous):

how did he do that a long __________________ line? =)

OpenStudy (anonymous):

over

OpenStudy (anonymous):

\[\cos(x)[2\sin(x)-1]\over[\cos ^{2}(x)+\sin ^{2}(x)+\sin ^{2}(x)-\cos ^{2}(x)]\]

OpenStudy (anonymous):

=\[\cos(x)[2\sin(x)-1]\over[2\sin ^{2}(x)-\sin(x)]\]

OpenStudy (anonymous):

= \[\cos(x)[2\sin(x)-1]\over \sin(x)[2\sin(x)-1]\]

OpenStudy (anonymous):

= cot(x)

OpenStudy (anonymous):

in the denominator of my first post there should be also a -sin(x)

OpenStudy (anonymous):

a ship started sailing S 42'35' W at the rate of 5kph. after 2 hours,ship B started at the same port going N 46'20' W at the rate of 7 kph.after how many hours will the second ship be exactly north of a ship A? can some answer this

OpenStudy (anonymous):

correct = cot(x)